if theres 180 grams of ci2 how many moles of FeCl3 can be produced?


1
Expert's answer
2022-03-22T16:42:03-0400

m(Cl2) = 180 g

n(FeCl3) - ?

Answer:

2Fe + 3Cl2 -> 2FeCl3

n(Cl2)= m/M = 180g/71g/mol = 2.53mol

n(FeCl3) = 3/2 n(Cl2) = 3.8 mol

Answer: 3.8 mol